# Robust Scene Coordinate Regression via Geometrically-Consistent Global Descriptors

### v1

@@ 5 Implementation details
Covisibility graph. Following R-Score [rscore], we construct the covisibility graph using pose-based overlap estimation [rau2020predicting]. For each training image, we sample random pixel coordinates and unproject them using uniformly sampled depth values to generate 3D point hypotheses. These 3D points are then projected onto the other camera view to compute pairwise overlap scores based on the fraction of points that fall within the image boundaries. An edge is added between two images to the covisibility graph if their overlap score exceeds a threshold of 0.2.
@@
Aggregator module. Our aggregator module processes dense DINO features to produce compact global descriptors. We employ the base version of DINOv2 (ViT-B/14) [dinov2], which generates 768-dimensional features. The aggregator consists of two main layers. The cluster layer reduces the dimensionality of DINO features from 768 to 128 dimensions using a trainable linear projection. The scoring layer then computes attention weights for each spatial location, enabling the aggregator to focus on the most relevant visual regions. The attention mechanism produces a weighted combination of the compressed features, resulting in an intermediate descriptor of 2304 dimensions. The intermediate descriptor is compressed to $256$ dimensions using a PCA layer [pca]. See Section 2 of the Supplementary Material for an ablation study on different dimensions.
@@
Aggregator training. We train the aggregator module for 10,000 iterations using a batch size of 64. The optimization employs AdamW [adam_kingma] with scene-specific learning rates: $3\times 10^{-3}$ for outdoor scenes and $1\times 10^{-3}$ for indoor scenes. The margin hyperparameter $\tau$ in Eq. 2 is set to $0.5$. The training takes less than an hour using a single NVIDIA H100 GPU. We train the aggregator without using any data augmentation, as the DINO features [dino] are already robust to various transformations.
@@
Coordinate regression settings. We follow the same configuration as R-Score [rscore]. Specifically, we use the DeDoDe keypoint encoder [edstedt2024dedode], which produces 128-dimensional local descriptors after being compressed using a PCA model. During testing, we adopt the multi-hypothesis strategy from R-Score [rscore], retrieving the top-$10$ hypotheses per query image. For this purpose, we use the SALAD [Izquierdo_CVPR_2024_SALAD] global descriptors with product quantization. We refer the reader to Table 3 for different results with other state-of-the-art global descriptors [ali2024boq, ali2023mixvpr, berton2025megaloc] and Section 5 of the Supplementary Material for other local descriptors.
@@
Coordinate regression training. We adopt the same optimization scheme as R-Score [rscore] to train the encoder module using a buffer of $128\times 10^{6}$ examples and a batch size of $320\times 10^{3}$. Instead of randomly sampling keypoints to populate the training buffer, we use heuristic sampling as described in FocusTune [focustune] using the SfM model if available. The training takes approximately 11 hours on average using a single NVIDIA H100 GPU. Finally, we use depth supervision and graph-based augmentation [rscore] for all of our experiments.
@@ 6 Evaluation
Figure 3: Median translation error. We plot the median translation error of the top-$k$ retrievals for all the training images. The results show that our learned global descriptors (in green) offer more relevant retrievals over the node2vec graph embeddings (in blue) of R-Score [rscore]. We also plot the ground-truth errors (in red) for reference.
Figure 4: Qualitative results. We visualize the top-$5$ retrievals for three images from the training set of the Aachen Day/Night dataset [sattler2018benchmarking]. For each query (left-most column), we show results using graph embeddings from R-Score [rscore] (top) and our learned global descriptors (bottom). On top of each retrieval, we plot the translation error and the overlap score between the retrievals and the query. Under heavy noise in the covisibility graph (evidenced by low overlap scores), the graph embeddings retrieve nearby but not exact structures, while our global descriptors retrieve nearby and relevant ones.
@@ 6.1 Quantitative evaluation
We evaluate our method on the Aachen day/night dataset [sattler2018benchmarking] and the Hyundai Department Store dataset [naverdatasets].
Aachen day/night dataset [sattler2018benchmarking]. Table 1 shows that our method enhances the accuracy of scene coordinate regression models [rscore, GLACE2024CVPR, brachmann2023accelerated, focustune]. Our method outperforms the state-of-the-art R-Score [rscore], achieving improvements of $1.1\%$ and $6.1\%$ for the $0.25\text{m}$ and $2^{\circ}$ threshold on daytime and nighttime images, respectively, and yielding an average gain of $2.2\%$ across all thresholds. Our method continues to close the gap to structure-based methods, such as hLoc [hloc] and DeViLoc [deviloc], achieving only $5.8\%$ and $5.4\%$ lower localization accuracy, respectively, while using just $0.67\%$ of hLoc’s memory.
Hyundai Department Store dataset [naverdatasets]. Table 2 shows that our method consistently outperforms R-Score [rscore] by $6.5\%$, $0.9\%$, and $4.1\%$ for floors 1F, 4F, and B1 respectively, for the $0.1\text{m}$ and $1^{\circ}$ threshold. Overall, we achieve a $2.5\%$ improvement in average localization accuracy while requiring only 7 MB additional memory. Compared to hLoc [hloc], our method further closes the gap to just $2.2\%$, using only $0.07\%$ of hLoc’s memory.
@@ 6.2 Qualitative evaluation
Figure 4 shows the top-$5$ retrievals for three randomly selected queries from the Aachen Day/Night training set [sattler2018benchmarking], comparing R-Score’s graph embeddings (top) with our learned global descriptors (bottom). The translation errors and the overlap scores are shown in the top-left and top-right corners, respectively. Our descriptors retain strong retrieval performance even under noisy geometrical constraints (evidenced by low overlap scores), due to their consistency with both the image space and the covisibility graph. In contrast, R-Score might retrieve irrelevant images whose visual content does not match the query image.
@@ 6.3 Ablation studies
Buffer sampling. Consistent with [focustune], we observe that using the SfM model to guide buffer sampling results in a more efficient training process and improved performance. As shown in Table 4, integrating FocusTune [focustune] yields nearly a $10\%$ improvement on the Aachen nighttime test set, highlighting the impact of effective buffer sampling for SCR models.
Global descriptors at query time. At test time, multiple hypotheses can be retrieved using any standard global descriptor model [DBLP:conf/cvpr/ArandjelovicGTP16, Izquierdo_CVPR_2024_SALAD, izquierdo2024close, ali2023mixvpr]. Table 3 presents the performance of several popular methods. Among them, SALAD [Izquierdo_CVPR_2024_SALAD] performs best on daytime queries, while BoQ [ali2024boq] excels at nighttime. Notably, even with NetVLAD [DBLP:conf/cvpr/ArandjelovicGTP16], our system achieves significant improvements over R-Score [rscore], gaining $1.3\%$ and $8.2\%$ at the $0.25\text{m}$ and $2^{\circ}$ threshold for day and night queries, respectively. These results highlight the superior quality of our learned global descriptors. We use SALAD [Izquierdo_CVPR_2024_SALAD] in all of our experiments.
Loss function for the aggregator module. Table 4 presents the impact of different loss functions used in the aggregator module. Our modified GCL (mGCL) consistently outperforms the original Generalized Contrastive Loss (GCL) [leyva2023gcl], yielding a $2.1\%$ accuracy gain on daytime images and a $3\%$ gain on nighttime images for the $0.25\text{m}$ and $2^{\circ}$ threshold on the Aachen dataset, highlighting the effectiveness of our mGCL in learning more discriminative global descriptors. The hyperparameter margin $\tau$ improves the accuracy of our system by $2-3\%$ compared to the default value of $0.5$ (see Section 1 of the Supplementary Material). We use mGCL with $\tau=0.5$ in all of our experiments.
@@
R-Score (with NetVLAD [DBLP:conf/cvpr/ArandjelovicGTP16]) 79.0 88.5 96.4 66.3 89.8 96.9
Ours (with NetVLAD [DBLP:conf/cvpr/ArandjelovicGTP16]) 79.9 91.0 97.2 73.5 90.8 96.9
Ours (with MegaLoc [berton2025megaloc]) 79.4 90.8 96.8 73.5 91.8 99.0
Ours (with MixVPR [ali2023mixvpr]) 79.9 90.8 97.2 73.5 89.8 96.9
Ours (with EigenPlaces [berton2023eigenplaces]) 79.5 90.3 96.0 70.4 91.8 96.9
Ours (with BoQ [ali2024boq]) 79.7 90.5 96.8 74.5 91.8 99.0
Ours (with SALAD [Izquierdo_CVPR_2024_SALAD]) 80.3 90.3 97.1 72.4 90.8 99.0
Table 3: Global descriptors at test time. To retrieve multiple hypotheses at test time, we can use any off-the-shelf global descriptors [DBLP:conf/cvpr/ArandjelovicGTP16, ali2023mixvpr, berton2023eigenplaces, ali2024boq, Izquierdo_CVPR_2024_SALAD]. We use SALAD [Izquierdo_CVPR_2024_SALAD] in all of our experiments.
@@
Table 4: Different components. Using FocusTune [focustune], better global descriptors, and our mGCL in Eq. 2 all further improve the performance by $8-14\%$ on the tightest threshold.
@@ 7 Conclusion
We introduce a novel aggregator module that learns geometrically-consistent global descriptors by enforcing dual consistency between visual similarity and geometrical connectivity via the covisibility graph. Unlike existing approaches that rely solely on the covisibility graph [rscore] or off-the-shelf models [GLACE2024CVPR], our method integrates both geometrical relationships and visual content to improve resilience to perceptual aliasing and noisy covisibility graphs.

### v2

@@ 5 Implementation details
Covisibility graph. Following R-Score [28], we construct the covisibility graph using pose-based overlap estimation [46]. For each training image, we sample random pixel coordinates and unproject them using uniformly sampled depth values to generate 3D point hypotheses. These 3D points are then projected onto the other camera view to compute pairwise overlap scores based on the fraction of points that fall within the image boundaries. An edge is added between two images to the covisibility graph if their overlap score exceeds a threshold of 0.2.
@@
Aggregator module. Our aggregator module processes dense DINO features to produce compact global descriptors. We employ the base version of DINOv2 (ViT-B/14) [43], which generates 768-dimensional features. The aggregator consists of two main layers. The cluster layer reduces the dimensionality of DINO features from 768 to 128 dimensions using a trainable linear projection. The scoring layer then computes attention weights for each spatial location, enabling the aggregator to focus on the most relevant visual regions. The attention mechanism produces a weighted combination of the compressed features, resulting in an intermediate descriptor of 2304 dimensions. The intermediate descriptor is compressed to $256$ dimensions using a PCA layer [22]. See Section 2 of the Supplementary Material for an ablation study on different dimensions.
@@
Aggregator training. We train the aggregator module for 10,000 iterations using a batch size of 64. The optimization employs AdamW [29] with scene-specific learning rates: $3\times 10^{-3}$ for outdoor scenes and $1\times 10^{-3}$ for indoor scenes. The margin hyperparameter $\tau$ in Eq. 2 is set to $0.5$. The training takes less than an hour using a single NVIDIA H100 GPU. We train the aggregator without using any data augmentation, as the DINO features [12] are already robust to various transformations.
@@
Coordinate regression settings. We follow the same configuration as R-Score [28]. Specifically, we use the DeDoDe keypoint encoder [19], which produces 128-dimensional local descriptors after being compressed using a PCA model. During testing, we adopt the multi-hypothesis strategy from R-Score [28], retrieving the top-$10$ hypotheses per query image. For this purpose, we use the SALAD [27] global descriptors with product quantization. We refer the reader to Table 3 for different results with other state-of-the-art global descriptors [2, 1, 5] and Section 5 of the Supplementary Material for other local descriptors.
@@
Coordinate regression training. We adopt the same optimization scheme as R-Score [28] to train the encoder module using a buffer of $128\times 10^{6}$ examples and a batch size of $320\times 10^{3}$. Instead of randomly sampling keypoints to populate the training buffer, we use heuristic sampling as described in FocusTune [41] using the SfM model if available. The training takes approximately 11 hours on average using a single NVIDIA H100 GPU. Finally, we use depth supervision and graph-based augmentation [28] for all of our experiments.
@@ 6 Evaluation
Figure 3: Median translation error. We plot the median translation error of the top-$k$ retrievals for all the training images. The results show that our learned global descriptors (in green) offer more relevant retrievals over the node2vec graph embeddings (in blue) of R-Score [28]. We also plot the ground-truth errors (in red) for reference.
Figure 4: Qualitative results. We visualize the top-$5$ retrievals for three images from the training set of the Aachen Day/Night dataset [54]. For each query (left-most column), we show results using graph embeddings from R-Score [28] (top) and our learned global descriptors (bottom). On top of each retrieval, we plot the translation error and the overlap score between the retrievals and the query. Under heavy noise in the covisibility graph (evidenced by low overlap scores), the graph embeddings retrieve nearby but not exact structures, while our global descriptors retrieve nearby and relevant ones.
@@ 6.1 Quantitative evaluation
We evaluate our method on the Aachen day/night dataset [54] and the Hyundai Department Store dataset [32].
Aachen day/night dataset [54]. Table 1 shows that our method enhances the accuracy of scene coordinate regression models [28, 60, 7, 41]. Our method outperforms the state-of-the-art R-Score [28], achieving improvements of $1.1\%$ and $6.1\%$ for the $0.25\text{m}$ and $2^{\circ}$ threshold on daytime and nighttime images, respectively, and yielding an average gain of $2.2\%$ across all thresholds. Our method continues to close the gap to structure-based methods, such as hLoc [48] and DeViLoc [20], achieving only $5.8\%$ and $5.4\%$ lower localization accuracy, respectively, while using just $0.67\%$ of hLoc’s memory.
Hyundai Department Store dataset [32]. Table 2 shows that our method consistently outperforms R-Score [28] by $6.5\%$, $0.9\%$, and $4.1\%$ for floors 1F, 4F, and B1 respectively, for the $0.1\text{m}$ and $1^{\circ}$ threshold. Overall, we achieve a $2.5\%$ improvement in average localization accuracy while requiring only 7 MB additional memory. Compared to hLoc [48], our method further closes the gap to just $2.2\%$, using only $0.07\%$ of hLoc’s memory.
@@ 6.2 Qualitative evaluation
Figure 4 shows the top-$5$ retrievals for three randomly selected queries from the Aachen Day/Night training set [54], comparing R-Score’s graph embeddings (top) with our learned global descriptors (bottom). The translation errors and the overlap scores are shown in the top-left and top-right corners, respectively. Our descriptors retain strong retrieval performance even under noisy geometrical constraints (evidenced by low overlap scores), due to their consistency with both the image space and the covisibility graph. In contrast, R-Score might retrieve irrelevant images whose visual content does not match the query image.
@@ 6.3 Ablation studies
Buffer sampling. Consistent with [41], we observe that using the SfM model to guide buffer sampling results in a more efficient training process and improved performance. As shown in Table 4, integrating FocusTune [41] yields nearly a $10\%$ improvement on the Aachen nighttime test set, highlighting the impact of effective buffer sampling for SCR models.
Global descriptors at query time. At test time, multiple hypotheses can be retrieved using any standard global descriptor model [3, 27, 26, 1]. Table 3 presents the performance of several popular methods. Among them, SALAD [27] performs best on daytime queries, while BoQ [2] excels at nighttime. Notably, even with NetVLAD [3], our system achieves significant improvements over R-Score [28], gaining $1.3\%$ and $8.2\%$ at the $0.25\text{m}$ and $2^{\circ}$ threshold for day and night queries, respectively. These results highlight the superior quality of our learned global descriptors. We use SALAD [27] in all of our experiments.
Loss function for the aggregator module. Table 4 presents the impact of different loss functions used in the aggregator module. Our modified GCL (mGCL) consistently outperforms the original Generalized Contrastive Loss (GCL) [34], yielding a $2.1\%$ accuracy gain on daytime images and a $3\%$ gain on nighttime images for the $0.25\text{m}$ and $2^{\circ}$ threshold on the Aachen dataset, highlighting the effectiveness of our mGCL in learning more discriminative global descriptors. The hyperparameter margin $\tau$ improves the accuracy of our system by $2-3\%$ compared to the default value of $0.5$ (see Section 1 of the Supplementary Material). We use mGCL with $\tau=0.5$ in all of our experiments.
@@
R-Score (with NetVLAD [3]) 79.0 88.5 96.4 66.3 89.8 96.9
Ours (with NetVLAD [3]) 79.9 91.0 97.2 73.5 90.8 96.9
Ours (with MegaLoc [5]) 79.4 90.8 96.8 73.5 91.8 99.0
Ours (with MixVPR [1]) 79.9 90.8 97.2 73.5 89.8 96.9
Ours (with EigenPlaces [6]) 79.5 90.3 96.0 70.4 91.8 96.9
Ours (with BoQ [2]) 79.7 90.5 96.8 74.5 91.8 99.0
Ours (with SALAD [27]) 80.3 90.3 97.1 72.4 90.8 99.0
Table 3: Global descriptors at test time. To retrieve multiple hypotheses at test time, we can use any off-the-shelf global descriptors [3, 1, 6, 2, 27]. We use SALAD [27] in all of our experiments.
@@
Table 4: Different components. Using FocusTune [41], better global descriptors, and our mGCL in Eq. 2 all further improve the performance by $8-14\%$ on the tightest threshold.
@@ 7 Conclusion
We introduce a novel aggregator module that learns geometrically-consistent global descriptors by enforcing dual consistency between visual similarity and geometrical connectivity via the covisibility graph. Unlike existing approaches that rely solely on the covisibility graph [28] or off-the-shelf models [60], our method integrates both geometrical relationships and visual content to improve resilience to perceptual aliasing and noisy covisibility graphs.